OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old. By how much does the replacement change the average age of the members of the orchestra?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

hw did you get this answer

OpenStudy (anonymous):

someone can explain this

OpenStudy (anonymous):

I don't know answer

jimthompson5910 (jim_thompson5910):

Let A = old average and B = new average The old average was found by dividing the sum of the ages S by 25 to get S/25 = A The new average is found by subtracting 32 from that sum S and adding in 30. Then you still divide by 25 (since you still have 25 people) (S-32+30)/25 = B (S-2)/25 = B S/25-2/25 = B A-2/25 = B B = A - 2/25 B = A - 0.08 So B = A - 0.08 which shows us that the average age decreased by 0.08 years